OpenStudy (brucebaner):

Katherine is having cloth napkins embroidered with her initials. The cost for the embroidery is $4.50 per napkin. Which graph represents the total cost for embroidering x napkins? X Y Z W

OpenStudy (brucebaner):

@AriPotta

OpenStudy (brucebaner):

help me @AriPotta

OpenStudy (brucebaner):

@chosenmatt

OpenStudy (brucebaner):

@KamiBug

OpenStudy (brucebaner):

@samson245

OpenStudy (brucebaner):

help me @AriPotta and @KamiBug

OpenStudy (aripotta):

so the line should have the equation y = 4.5x. find 2 points on the line by substituting 1 and 2 for x and finding y. y = 4.5(1) = 4.5 (1, 4.5) y = 4.5(2) = 9 (2, 9) so look for the choice with points (1, 4.5) and (2, 9)

OpenStudy (brucebaner):

so its Z

OpenStudy (aripotta):

yep

OpenStudy (brucebaner):

thx for helping me @AriPotta

OpenStudy (aripotta):

no problemo :)